OpenStudy (anonymous):

What is the solution of the equation? 36x2 = 100

OpenStudy (anonymous):

do you mean 36 x 2 = 100 or 36x^2=100?

OpenStudy (anonymous):

36x^2=100? i think the ansers is -9 over 25,and9over25

OpenStudy (anonymous):

my answer is 1,67

OpenStudy (anonymous):

tahst not one of the ansers

OpenStudy (anonymous):

the answer is find x?

OpenStudy (anonymous):

i i got i never mind

OpenStudy (anonymous):

36x^2 = 100 x^2 = 100/36 x = +/- (10/6) 1. What is the solution of the equation? 36x2 = 100 (1 point) (0 pts) –9/25, 9/25 (0 pts) –3/5, 3/5 (0 pts) –25/9, 25/9 (1 pt) –5/3, 5/3 1 /1 point